# Compact Binary Formation in Open Star Clusters II: Difficulty of Gaia NS formation in low-mass star clusters

### v1

@@ Abstract
Gaia mission offers opportunities to search for compact binaries not involved in binary interactions (hereafter inert compact binaries), and results in the discoveries of binaries containing one black hole (BH) or one neutron star (NS), called “Gaia BHs” and “Gaia NSs”, respectively. We have assessed if Gaia BHs and NSs can be formed in open clusters through dynamical interactions. In order to obtain a large number of inert compact binaries similar to Gaia BHs and NSs, we have performed gravitational $N$-body simulations for a large number of open clusters whose total mass is $1.2\times 10^{8}M_{\odot}$. These clusters have various masses, metallicities, densities, and binary fractions. We have found that open clusters form Gaia BHs ($10^{-6}$-$10^{-5}M_{\odot}^{-1}$) much more efficiently than Gaia NSs ($\lesssim 10^{-7}M_{\odot}^{-1}$) for any cluster parameters. This is quite inconsistent with observational results, because the reported numbers of Gaia BHs and NSs are $2$ and $\sim 20$, respectively. Additionally, we have switched off NS natal kicks for $10^{4}$ open clusters each weighing $10^{3}M_{\odot}$ in order to retain a large number of NSs in open clusters. Then, open clusters form inert NS binaries originating from primordial binaries rather than formed through dynamical interactions. This means that Gaia NSs are formed dominantly on isolated fields, not in open clusters, if there is no NS natal kick. We have concluded that Gaia BHs can be dominantly formed in open clusters, however Gaia NSs cannot.
@@ 1 Introduction
Very recently, inactive compact objects have started to be discovered. Mainly, they are binary members, because they are found through binary motions. They are called “non-interacting”, “detached”, “dormant”, or “inert” compact binaries. Hereafter, we unify the notation with inert compact binaries throughout this paper. Such inert compact binaries have been discovered by spectroscopic observations (Giesers et al., 2018; Shenar et al., 2022), and by astrometric observations (El-Badry et al., 2023b, c; Tanikawa et al., 2023; Chakrabarti et al., 2023) with the help of Gaia DR3 (Gaia Collaboration et al. 2023a, b; see review by El-Badry 2024). Gaia DR3 seems to contain more inert compact binaries according to several candidate lists (Andrews et al., 2022; Shahaf et al., 2023b; Shahaf et al., 2023a; Jayasinghe et al., 2023; Rowan et al., 2024). Although other inert compact binaries have been reported (Liu et al., 2019; Thompson et al., 2019; Rivinius et al., 2020; Jayasinghe et al., 2021, 2022; Lennon et al., 2022; Saracino et al., 2022), their discoveries have been called into question (Abdul-Masih et al., 2020; Bodensteiner et al., 2020; El-Badry & Quataert, 2020, 2021; van den Heuvel & Tauris, 2020; El-Badry & Burdge, 2022; El-Badry et al., 2022a; El-Badry et al., 2022b). In addition to inert compact binaries, several single BHs/NSs have been detected by microlensing observations (Lam et al., 2022; Sahu et al., 2022; Howil et al., 2024).
@@
Gaia BH1 and BH2 (hereafter Gaia BHs), which are inert BH binaries discovered from Gaia DR3, and named by El-Badry et al. (2023b, c), have two impacts. First, Gaia BHs have the longest orbital periods among BH binaries discovered so far. In particular, Gaia BH2 breaks the record of BH binary periods by 10 times. These discoveries give the impression that Gaia’s astrometric observations open up a new parameter space of BH binaries. Second, Gaia BHs are hard to form through isolated binary evolution. If Gaia BHs are formed through isolated binary evolution, common envelope ejection should be 10 times less efficient than expected (Chawla et al. 2022; El-Badry et al. 2023b; Shikauchi et al. 2023; but see Kotko et al. 2024). However, a recent study has not supported such inefficient common envelope ejection for stars massive enough to form BHs (Hirai & Mandel, 2022). Triple stars (El-Badry et al., 2023b, c; Generozov & Perets, 2023) and open clusters (Di Carlo et al., 2023; Rastello et al., 2023; Tanikawa et al., 2024) have been suggested as formation sites of Gaia BHs. Because Gaia BHs are the Galactic disk components, it is difficult to suppose that they are formed in globular clusters and galactic centers.
@@
El-Badry et al. (2024) have discovered Gaia NS1 which is most likely to be the first inert NS binary from Gaia DR3 (Gaia Collaboration et al., 2023b). The NS candidate mass ($\sim 1.9M_{\odot}$) is higher than other inert NS binary candidates which could be a massive white dwarfs (WDs) as described in the literature (Geier et al., 2023; Zheng et al., 2023). Gaia NS1 is hard to form through isolated binary evolution without inefficient common envelope ejection, similarly to Gaia BHs. El-Badry et al. (2024) have also mentioned that they have found about 20 more inert NS binaries, called “Gaia NSs” hereafter. This should imply that Gaia NSs are formed more efficiently than Gaia BHs for the following two reasons. First, the number of Gaia NSs discovered is larger than that of Gaia BHs discovered. Second, Gaia NSs is harder to detect than Gaia BHs, because NSs indicate smaller astrometric signals than BHs; NSs swing around their companions less strongly than BHs due to their smaller masses. Another importance of Gaia NS1 is its eccentricity. The eccentricity is $\sim 0.1$ (El-Badry et al., 2024), while Gaia BH1 and BH2 have $\sim 0.5$ (El-Badry et al., 2023b, c). The eccentricity of Gaia NS1 may be too small to be formed through dynamical interactions.
@@ 2 Method
NS population retained in open clusters largely depends on NS natal kicks. Here, we remark NS natal kicks in our model. NSs receive kick velocities at their births because of asymmetric supernova explosions. The kick velocities are quite different between NSs born from electron capture (EC) and core collapse (CC) supernovae (SNe). In our model, a star causes a CCSN when its carbon-oxygen (CO) core exceeds both of the Chandrasekhar mass limit ($1.44M_{\odot}$ here) and $0.773M_{\rm c,he}-0.35M_{\odot}$, where $M_{\rm c,he}$ is the He core mass of the star. NSs left behind CCSNe have kick velocity distribution given by a single Maxwellian with $\sigma_{\rm k,ccsn}=265$ km s-1 (Hobbs et al., 2005). If the NSs acquires on fallback mass, the kick velocities are reduced by $1-f_{\rm fallback}$, where $f_{\rm fallback}$ is the fraction of the fallback mass (Fryer et al., 2012). This prescription is also applied to BH natal kicks. ECSNe (Miyaji et al., 1980; Nomoto, 1984, 1987; van den Heuvel, 2007) are thought to raise much smaller kick velocities than CCSNe because of slightly asymmetric explosion (Jones et al., 2013; Gessner & Janka, 2018). In our model, a star generates an ECSN when it does not cause a CCSN, and its CO core mass is $\geq 1.6M_{\odot}$. Their natal kick velocities are expressed by a single Maxwellian with $3$ km s-1 (Podsiadlowski et al., 2004; Gessner & Janka, 2018). About $1$ of NSs are formed from ECSNe under the initial mass function we adopt.
@@
Additionally, we make the no-kick model in which no CCSNe generate kick velocities, in order to investigate effects of CCSN natal kicks. We set the metallicity to $Z=0.002$, because the visible star of Gaia NS1 has small metallicity, $[{\rm Fe/H}]=-1.23$ (El-Badry et al., 2024).
We generate a large number of open clusters with different realizations, so that each model contains $10^{7}M_{\odot}$ open clusters in total. The total mass of simulated clusters is $1.3\times 10^{8}M_{\odot}$. We can expect that we obtain $\sim 100$ Gaia BHs from each cluster model, because the formation efficiency of Gaia BHs is $\sim 10^{-6}M_{\odot}^{-1}$ (Rastello et al., 2023; Di Carlo et al., 2023; Tanikawa et al., 2024). This enables us to make comparison between the formation efficiencies of Gaia BHs and NSs.
@@
    $10^{2}\leq P/{\rm day}\leq 10^{4}$, where $P$ is a binary orbital period. This criterion may look loose, since the orbital periods of Gaia BH1, BH2, and NS1 are $180$, $1300$, and $730$ days, respectively. Thus, we also investigate the case for $10^{2}\leq P/{\rm day}\leq 10^{3}$.
@@
    $m_{2}\leq 1.1M_{\odot}$, where $m_{2}$ is a companion mass. This is because Gaia BH1, BH2, and NS1 have light companion stars: $0.93$, $1.07$, and $0.79M_{\odot}$, respectively.
@@
    Located outside of open clusters. None of Gaia BH1, BH2 and NS1 are located in open clusters.
Note that we do not put any constraints on orbital eccentricities in the above criteria. On the other hand, Gaia BHs and NS1 are characterized by moderately high eccentricities ($\sim 0.5$) and small (but non-zero) eccentricity ($\sim 0.1$), respectively. Compact binaries we pick up may not be necessarily identified as Gaia BHs and NSs. Thus, these compact binaries are hereafter called “inert compact binaries”. They are also called “inert BH binaries (inert BHBs)” and “inert NS binaries (inert NSBs)” if they contain BHs and NSs, respectively.
@@ 3.1 Standard CCSN natal kicks
Figure 1 shows the formation efficiencies of inert BHBs and NSBs in all our cluster models other than the no-kick model. Inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ are formed at a formation efficiency of $10^{-6}$-$10^{-5}M_{\odot}^{-1}$ except for the $f_{\rm b,<5M_{\odot}}=0$ model. This is consistent with previous results (Di Carlo et al., 2023; Rastello et al., 2023; Tanikawa et al., 2024). The formation efficiencies slowly increase with metallicity decreasing, initial cluster mass increasing, and initial half-mass density increasing. This is because more metal-poor, larger-mass, and higher-density clusters are harder to be disrupted in the Galactic tidal field, and thus these clusters have longer time to form inert BHBs. The formation efficiency in the $f_{\rm b,<5M_{\odot}}=0$ model is much smaller than those in the other models for the following reason. Generally, an inert BHB is formed, such that a single BH captures one star in a binary. Since the number of binaries in the $f_{\rm b,<5M_{\odot}}=0$ model is much smaller than those in the other models, inert BHBs are harder to form in the $f_{\rm b,<5M_{\odot}}=0$ model. Nevertheless, it is worth remarking that inert BHBs are formed at an efficiency of $\sim 10^{-7}M_{\odot}^{-1}$ even in the $f_{\rm b,<5M_{\odot}}=0$ model. They would be formed through triple-single encounters. This indicates that inert BHBs are robustly formed in open clusters, even if an initial binary fraction is extremely low. This is consistent with the results of Tanikawa et al. (2024).
@@
Figure 2: Cumulative eccentricity distribution of inert BHBs and NSBs. Solid and dashed curves indicate inert BHBs with binary orbital periods of $10^{2}$-$10^{4}$ and $10^{2}$-$10^{3}$ days, respectively. The dot dashed curve indicates inert NSBs in all the models. The dotted curve shows the thermal distribution.
Figure 2 shows the orbital eccentricity distribution of inert BHBs and NSBs. We put together inert BHBs and NSBs among all the models other than the no-kick model. We have to put together inert NSBs, because the number of inert NSBs is too small to statistically investigate the eccentricity distribution of inert NSBs for each model. We adopt the Kolmogorov-Smirnov (K-S) test to compare these distribution with the thermal distribution (Heggie, 1975) and Gaia BHs and NSs. Note that Gaia BH1, BH2 and NS1 have orbital eccentricities of $\sim 0.451$ (El-Badry et al., 2023b), $0.518$ (El-Badry et al., 2023c), and $0.124$ (El-Badry et al., 2024), respectively. Table 2 shows the P-values of the K-S test. The eccentricity distribution of inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ is clearly more circular than the thermal distribution, since the P-value is $5.6\times 10^{-5}$. This is possibly because eccentric BHBs are circularized by tidal interactions. This is also seen in the eccentricity-period diagram of binaries in Gaia DR3 (e.g. Gaia Collaboration et al., 2023b). On the other hand, the eccentricity distribution of inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{3}$ is not significantly different from the thermal distribution; the P-value is $9.3\times 10^{-2}$. This may be due to their small number.
@@
Inert BHB ($10^{2}$-$10^{4}$ day) 556 Gaia BHs $2.4\times 10^{-1}$
Inert BHB ($10^{2}$-$10^{3}$ day) 97 Gaia BHs $2.7\times 10^{-1}$
@@
Inert BHBs and NSBs have the eccentricity distributions which are not inconsistent with Gaia BHs and NSs, respectively, regardless of how to choose these BHBs and NSBs. The P-values are more than $0.1$. This may be due to the small numbers; the number of inert NSBs are only 6, and the numbers of Gaia BHs and NSs are 2 and 1, respectively. It seems that the eccentricity distributions are not currently helpful to assess if inert BHBs and NSBs are identified as Gaia BHs and NSs, respectively.

### v2

@@ Abstract
Gaia mission offers opportunities to search for compact binaries not involved in binary interactions (hereafter inert compact binaries), and results in the discoveries of binaries containing one black hole (BH) or one neutron star (NS), called “Gaia BHs” and “Gaia NSs”, respectively. We have assessed if Gaia BHs and NSs can be formed in open clusters through dynamical interactions. In order to obtain a large number of inert compact binaries similar to Gaia BHs and NSs, we have performed gravitational $N$-body simulations for a large number of open clusters whose total mass is $1.2\times 10^{8}M_{\odot}$. These clusters have various masses, metallicities, densities, and binary fractions. We have found that open clusters form Gaia BHs ($10^{-6}$-$10^{-5}M_{\odot}^{-1}$) much more efficiently than Gaia NSs ($\lesssim 10^{-7}M_{\odot}^{-1}$) for any cluster parameters. This is quite inconsistent with observational results, because the reported numbers of Gaia BHs and NSs are $3$ and $21$, respectively. Additionally, we have switched off NS natal kicks for $10^{4}$ open clusters each weighing $10^{3}M_{\odot}$ in order to retain a large number of NSs in open clusters. Then, open clusters form inert NS binaries originating from primordial binaries rather than formed through dynamical interactions. This means that Gaia NSs are formed dominantly on isolated fields, not in open clusters, if there is no NS natal kick. We have concluded that Gaia BHs can be dominantly formed in open clusters, however Gaia NSs cannot.
@@ 1 Introduction
Very recently, inactive compact objects have started to be discovered. Mainly, they are binary members, because they are found through binary motions. They are called “non-interacting”, “detached”, “dormant”, or “inert” compact binaries. Hereafter, we unify the notation with inert compact binaries throughout this paper. Such inert compact binaries have been discovered by spectroscopic observations (Giesers et al., 2018; Shenar et al., 2022), and by astrometric observations (El-Badry et al., 2023b, c; Tanikawa et al., 2023; Chakrabarti et al., 2023) with the help of Gaia DR3 (Gaia Collaboration et al. 2023a, b; see review by El-Badry 2024a). Gaia DR3 seems to contain more inert compact binaries according to several candidate lists (Andrews et al., 2022; Shahaf et al., 2023b; Shahaf et al., 2023a; Jayasinghe et al., 2023; Rowan et al., 2024). Although other inert compact binaries have been reported (Liu et al., 2019; Thompson et al., 2019; Rivinius et al., 2020; Jayasinghe et al., 2021, 2022; Lennon et al., 2022; Saracino et al., 2022), their discoveries have been called into question (Abdul-Masih et al., 2020; Bodensteiner et al., 2020; El-Badry & Quataert, 2020, 2021; van den Heuvel & Tauris, 2020; El-Badry & Burdge, 2022; El-Badry et al., 2022a; El-Badry et al., 2022b). In addition to inert compact binaries, several single BHs/NSs have been detected by microlensing observations (Lam et al., 2022; Sahu et al., 2022; Howil et al., 2024).
@@
Gaia BH1 and BH2 (hereafter Gaia BHs), which are inert BH binaries discovered from Gaia DR3, and named by El-Badry et al. (2023b, c), have two impacts. First, Gaia BHs have the longest orbital periods among BH binaries discovered so far. In particular, Gaia BH2 breaks the record of BH binary periods by 10 times. These discoveries give the impression that Gaia’s astrometric observations open up a new parameter space of BH binaries. Second, Gaia BHs are hard to form through isolated binary evolution. If Gaia BHs are formed through isolated binary evolution, common envelope ejection should be 10 times more efficient than expected (Chawla et al. 2022; El-Badry et al. 2023b; Shikauchi et al. 2023; but see Kotko et al. 2024). However, a recent study has not supported such efficient common envelope ejection for stars massive enough to form BHs (Hirai & Mandel, 2022). Triple stars (El-Badry et al., 2023b, c; Generozov & Perets, 2023) and open clusters (Shikauchi et al., 2020; Rastello et al., 2023; Tanikawa et al., 2024; Di Carlo et al., 2024) have been suggested as formation sites of Gaia BHs. Because Gaia BHs are the Galactic disk components, it is difficult to suppose that they are formed in globular clusters and galactic centers.
During the review process of this paper, Gaia BH3 has been reported by Gaia Collaboration et al. (2024). It could belong to the ED-2 stream, which is possibly a tidally disrupted globular cluster (Balbinot et al., 2024). Thus, it would be formed through dynamical interactions in the globular cluster (Marín Pina et al., 2024), which is similar to Gaia BH formation in open clusters (Shikauchi et al., 2020; Rastello et al., 2023; Tanikawa et al., 2024; Di Carlo et al., 2024). However, Gaia BH3 could be formed from isolated binaries (El-Badry, 2024b; Iorio et al., 2024), since it is a very metal-poor binary, [Fe/H] $=-2.56$ (Gaia Collaboration et al., 2024). The origin of Gaia BHs is an open question.
@@
El-Badry et al. (2024b) have discovered Gaia NS1 which is most likely to be the first inert NS binary from Gaia DR3 (Gaia Collaboration et al., 2023b). The NS candidate mass ($\sim 1.9M_{\odot}$) is higher than other inert NS binary candidates which could be a massive white dwarfs (WDs) as described in the literature (Geier et al., 2023; Zheng et al., 2023). Gaia NS1 is hard to form through isolated binary evolution without efficient common envelope ejection, similarly to Gaia BHs. El-Badry et al. (2024b) have also mentioned that they have found about 20 more inert NS binaries, called “Gaia NSs” hereafter. El-Badry et al. (2024a) have published their detail information during the review process of this paper. This should imply that Gaia NSs are formed more efficiently than Gaia BHs for the following two reasons. First, the number of Gaia NSs discovered is larger than that of Gaia BHs discovered. Second, Gaia NSs is harder to detect than Gaia BHs, because NSs indicate smaller astrometric signals than BHs; NSs swing around their companions less strongly than BHs due to their smaller masses. Another importance of Gaia NS1 is its eccentricity. The eccentricity is $\sim 0.1$ (El-Badry et al., 2024b), while Gaia BH1 and BH2 have $\sim 0.5$ (El-Badry et al., 2023b, c). The eccentricity of Gaia NS1 may be too small to be formed through dynamical interactions. We also look into the eccentricities of all the Gaia NSs.
@@ 2 Method
NS population retained in open clusters largely depends on NS natal kicks. Here, we remark on NS natal kicks in our model. NSs receive kick velocities at their births because of asymmetric supernova explosions. The kick velocities are quite different between NSs born from electron capture (EC) and core collapse (CC) supernovae (SNe). In our model, a star causes a CCSN when its carbon-oxygen (CO) core exceeds both of the Chandrasekhar mass limit ($1.44M_{\odot}$ here) and $0.773M_{\rm c,he}-0.35M_{\odot}$, where $M_{\rm c,he}$ is the He core mass of the star. NSs left behind CCSNe have kick velocity distribution given by a single Maxwellian with $\sigma_{\rm k,ccsn}=265$ km s-1 (Hobbs et al., 2005). If the NSs acquires on fallback mass, the kick velocities are reduced by $1-f_{\rm fallback}$, where $f_{\rm fallback}$ is the fraction of the fallback mass (Fryer et al., 2012). This prescription is also applied to BH natal kicks. ECSNe (Miyaji et al., 1980; Nomoto, 1984, 1987; van den Heuvel, 2007) are thought to raise much smaller kick velocities than CCSNe because of slightly asymmetric explosion (Jones et al., 2013; Gessner & Janka, 2018). In our model, a star generates an ECSN when it does not cause a CCSN, and its CO core mass is $\geq 1.6M_{\odot}$. Their natal kick velocities are expressed by a single Maxwellian with $3$ km s-1 (Podsiadlowski et al., 2004; Gessner & Janka, 2018). About $1$ of NSs are formed from ECSNe under the initial mass function we adopt.
@@
Additionally, we make the no-kick model in which no CCSNe generate kick velocities, in order to investigate effects of CCSN natal kicks. We set the metallicity to $Z=0.002$, because the visible star of Gaia NS1 has small metallicity, $[{\rm Fe/H}]=-1.23$ (El-Badry et al., 2024b).
We generate a large number of open clusters with different realizations, so that each model contains $10^{7}M_{\odot}$ open clusters in total. The total mass of simulated clusters is $1.3\times 10^{8}M_{\odot}$. We can expect that we obtain $\sim 100$ Gaia BHs from each cluster model, because the formation efficiency of Gaia BHs is $\sim 10^{-6}M_{\odot}^{-1}$ (Rastello et al., 2023; Tanikawa et al., 2024; Di Carlo et al., 2024). This enables us to make comparison between the formation efficiencies of Gaia BHs and NSs.
@@
    $10^{2}\leq P/{\rm day}\leq 10^{4}$, where $P$ is a binary orbital period. This criterion may look loose, since the orbital periods of Gaia BHs and Gaia NSs are $100$-$4000$ days, and $100-1000$ days, respectively. Thus, we also investigate the case for $10^{2}\leq P/{\rm day}\leq 10^{3}$.
@@
    $m_{2}\leq 1.1M_{\odot}$, where $m_{2}$ is a companion mass. This is because Gaia BH1, BH2, and NS1 have light companion stars: $0.93$, $1.07$, and $0.79M_{\odot}$, respectively. This is also true for Gaia BH3 and a large fraction of Gaia NSs.
@@
    Located outside of open clusters. None of Gaia BHs and Gaia NSs are located in open clusters.
Note that we do not put any constraints on orbital eccentricities in the above criteria. On the other hand, Gaia BHs and NS1 are characterized by moderately high eccentricities ($0.5$-$0.7$) and small (but non-zero) eccentricity ($\sim 0.1$), respectively. Note that the eccentricities of Gaia NSs range from $0.1$ to $0.8$. Compact binaries we pick up may not be necessarily identified as Gaia BHs and NSs. Thus, these compact binaries are hereafter called “inert compact binaries”. They are also called “inert BH binaries (inert BHBs)” and “inert NS binaries (inert NSBs)” if they contain BHs and NSs, respectively.
@@ 3.1 Standard CCSN natal kicks
Figure 1 shows the formation efficiencies of inert BHBs and NSBs in all our cluster models other than the no-kick model. Inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ are formed at a formation efficiency of $10^{-6}$-$10^{-5}M_{\odot}^{-1}$ except for the $f_{\rm b,<5M_{\odot}}=0$ model. This is consistent with previous results (Rastello et al., 2023; Tanikawa et al., 2024; Di Carlo et al., 2024). The formation efficiencies slowly increase with metallicity decreasing, initial cluster mass increasing, and initial half-mass density increasing. The reason for the dependence on initial cluster mass and half-mass density is that larger-mass, and higher-density clusters are harder to be disrupted in the Galactic tidal field, and thus these clusters have longer time to form inert BHBs. We describe the reason for the dependence on metallicity later. The formation efficiency in the $f_{\rm b,<5M_{\odot}}=0$ model is much smaller than those in the other models for the following reason. Generally, an inert BHB is formed, such that a single BH captures one star in a binary. Since the number of binaries in the $f_{\rm b,<5M_{\odot}}=0$ model is much smaller than those in the other models, inert BHBs are harder to form in the $f_{\rm b,<5M_{\odot}}=0$ model. Nevertheless, it is worth remarking that inert BHBs are formed at an efficiency of $\sim 10^{-7}M_{\odot}^{-1}$ even in the $f_{\rm b,<5M_{\odot}}=0$ model. They would be formed through triple-single encounters. This indicates that inert BHBs are robustly formed in open clusters, even if an initial binary fraction is extremely low. This is consistent with the results of Tanikawa et al. (2024).
@@
The inert BHB formation efficiencies weakly increase with metallicities decreasing. The formation efficiency in the $Z=0.0002$ model is 2 and 5 times than in the $Z=0.02$ model for inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ and $10^{2}\leq P/{\rm day}\leq 10^{3}$, respectively. We look into the reason of the dependence on metallicity. This is because the dependence may be a key to elucidate the origin of Gaia BHs as discussed in El-Badry (2024b). Single BH formation efficiencies are not sensitive to metallicities: $2.7\times 10^{-3}$, $2.8\times 10^{-3}$, $2.9\times 10^{-3}$, $3.2\times 10^{-3}$, and $3.0\times 10^{-3}M_{\odot}^{-1}$ for $Z=0.02$, $0.01$, $0.005$, $0.002$, and $0.0002$ models, respectively. Metallicities do not affect cluster mass evolution: $\sim 700M_{\odot}$, $\sim 200M_{\odot}$ and $\lesssim 1M_{\odot}$ at the clusters’ ages of $40$, $400$, and $1000$ Myrs, respectively, for all the metallicities. Thus, single BH formation efficiency and cluster mass evolution are not the reason for the dependence of the inert BHB formation efficiencies on metallicities. On the other hand, the average single BH masses are quite different among metallicities: $7.9$, $11.7$, $15.7$, $17.4$, and $20.3M_{\odot}$ for $Z=0.02$, $0.01$, $0.005$, $0.002$, and $0.0002$ models, respectively. Since more massive BHs can capture other stars more easily, single BH masses should strongly affect the dependence of the inert BHB formation efficiencies.
As seen in Figure 1, the formation efficiencies of inert BHBs weakly increase with increasing cluster mass. On the other hand, Figure 2 in Marín Pina et al. (2024) has shown that the formation efficiencies decrease with increasing cluster mass. This difference would come from two points. First, inert BHB formation is inhibited due to more pronounced dynamical heating by larger BH populations in more massive clusters, as pointed out by Kremer et al. (2018) and Marín Pina et al. (2024). The second point is as follows. For clusters with $\lesssim 10^{3}M_{\odot}$, two-body relaxation (or thermal) timescale (Spitzer, 1987) is sufficiently shorter than the Hubble time. Thus, BHs can interact with other stars for a long time with respect to clusters’ thermal timescales. However, their interactions are prevented by tidal disruption of clusters. Since more massive clusters are more robust against the Galactic tidal field, they can form inert BHBs more efficiently. For clusters with $\gtrsim 10^{5}M_{\odot}$, two-body relaxation time is comparable to or longer than the Hubble time. From the view point of clusters’ thermal timescales, BHs have a shorter time to interact with other stars in more massive clusters. Note that the two-body relaxation time increases with increasing cluster mass. In fact, the number of BHs decreases more slowly with increasing cluster mass as seen in Figure 3 of Kremer et al. (2020), which have published the CMC models used by Marín Pina et al. (2024). This clearly indicates that clusters’ thermal ages are younger and BHs have less chances to interact with other stars in more massive stars. About $10^{4}M_{\odot}$ clusters may form inert BHs most efficiently, because they are hard to be tidally disrupted, and have a sufficiently shorter two-body relaxation time than the Hubble time.
@@
El-Badry et al. (2024b, a) have not fully ruled out the possibility that some of Gaia NSs are triple systems; for example NSs are actually double WDs. This motivates us to investigate triple and quadruple systems formed in all our cluster models. Such systems include an inner MS binary with a single NS as well as an inner double WDs with a single MS star. We do not find such a triple system. Instead, we find only one triple system which is potentially mistook for an inert NSB from the $Z=0.0002$ model. Its inner binary consists of a $1.28M_{\odot}$ NS and $0.022M_{\odot}$ HeWD, and its period and eccentricity are $32$ minutes and $0.023$, respectively. This inner binary is formed from a primordial binary. The outer binary has the period and eccentricity of $560$ days and $0.76$, respectively, and the other outer component is a $0.95M_{\odot}$ MS. The inner binary is in the process of mass transfer from the HeWD to the NS. This mass transfer is stable and long-lasting. This is the reason why the inner binary can capture the third star. Because of the mass transfer, the triple system should be observed as an X-ray binary, not as an inert NSB. Thus, we exclude the triple system from our list of inert NSBs. Even if we add it to our list, the formation efficiency of inert NSBs is still much smaller than that of inert BHBs.
Figure 2: Cumulative eccentricity distribution of inert BHBs and NSBs. Solid and dashed curves indicate inert BHBs with binary orbital periods of $10^{2}$-$10^{4}$ and $10^{2}$-$10^{3}$ days, respectively. The dot dashed curve indicates inert NSBs in all the models. The thick solid curve indicates all the Gaia NSs. The dotted curve shows the thermal distribution.
Figure 2 shows the orbital eccentricity distribution of inert BHBs and NSBs. We put together inert BHBs and NSBs among all the models other than the no-kick model. We have to put together inert NSBs, because the number of inert NSBs is too small to statistically investigate the eccentricity distribution of inert NSBs for each model. We adopt the Kolmogorov-Smirnov (K-S) test to compare these distribution with the thermal distribution (Heggie, 1975) and Gaia BHs and NSs. Note that Gaia BH1, BH2 and NS1 have orbital eccentricities of $\sim 0.451$ (El-Badry et al., 2023b), $0.518$ (El-Badry et al., 2023c), and $0.124$ (El-Badry et al., 2024b), respectively. After this paper was submitted, El-Badry et al. (2024a) have also published the orbital parameters of all Gaia NSs. We compare inert NSBs not only with Gaia NS1, but also with all the Gaia NSs. Table 2 shows the P-values of the K-S test. The eccentricity distribution of inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{4}$ is clearly more circular than the thermal distribution, since the P-value is $5.6\times 10^{-5}$. This is possibly because eccentric BHBs are circularized by tidal interactions. This is also seen in the eccentricity-period diagram of binaries in Gaia DR3 (e.g. Gaia Collaboration et al., 2023b). On the other hand, the eccentricity distribution of inert BHBs with $10^{2}\leq P/{\rm day}\leq 10^{3}$ is not significantly different from the thermal distribution; the P-value is $9.3\times 10^{-2}$. This may be due to their small number.
@@
Inert BHB ($10^{2}$-$10^{4}$ day) 556 Gaia BHs $6.4\times 10^{-1}$
Inert BHB ($10^{2}$-$10^{3}$ day) 97 Gaia BHs $6.9\times 10^{-1}$
@@
Inert NSB ($10^{2}$-$10^{4}$ day) 6 All Gaia NSs $1.9\times 10^{-4}$
Inert NSB selected 4 All Gaia NSs $2.4\times 10^{-3}$
Inert NSB (no-kick) 34 All Gaia NSs $1.0\times 10^{-5}$
@@
Inert BHBs and NSBs have the eccentricity distributions which are not inconsistent with Gaia BHs and Gaia NS1, respectively, regardless of how to choose these BHBs and NSBs. The P-values are more than $0.1$. This may be due to the small numbers; the number of inert NSBs are only 6, and the numbers of Gaia BHs and Gaia NS1 are 2 and 1, respectively. In fact, when we compare our inert NSBs with all the Gaia NSs, the P-values are quite small: $1.9\times 10^{-4}$ for all the inert NSBs, and $2.4\times 10^{-3}$ for the inert NSBs selected. Owing to the publication of Gaia NSs’ eccentricities, we can find that the eccentricity distribution of our inert NSBs is different from that of Gaia NSs. On the other hand, It seems that the eccentricity distribution of Gaia BHs is not currently helpful to assess if inert BHBs are identified as Gaia BHs.
@@
If we reduce the minimum binary mass ratio, another inert NSB formation channel (hereafter called primordial binary or PB channel) similar to the CCSN channel may appear in primordial binaries, and in isolated binaries. The difference between the two channels is how to form progenitors of inert NSBs. The progenitors are formed through dynamical interactions in the CCSN channel, while they are present from the initial time in the PB channel. The PB channel increases the formation efficiency of inert NSBs (possibly BHBs) not only in open clusters, but also on isolated fields. In this case, the importance of open clusters would diminish. This is because the number of channels increases from 3 to 4 in open clusters, while it increases from 0 to 1 on isolated fields.